# Adaptive Vision-Guided Robotic Arm Control for Precision Pruning in Dynamic Orchard Environments

### v1

@@ Abstract
This study presents a vision-guided robotic control system for automated fruit tree pruning applications. Traditional agricultural practices rely on labor-intensive tasks and processes that lack scalability and efficiency, creating a pressing need for automation research to address growing demands for higher crop yields, scalable operations, and reduced manual labor. To this end, this paper proposes a novel algorithm for robust and automated fruit pruning in dense orchards. The proposed algorithm utilizes CoTracker, that is designed to track 2D feature points in video sequences with significant robustness and accuracy, while leveraging joint attention mechanisms to account for inter-point dependencies, enabling robust and precise tracking under challenging and sophisticated conditions. To validate the efficacy of CoTracker, a Universal Robots manipulator UR5e is employed in a Gazebo simulation environment mounted on ClearPath Robotics Warthog robot featuring an Intel RealSense D435 camera. The system achieved a 93% success rate in pruning trials and with an average end trajectory error of 0.23 mm. The vision controller demonstrated robust performance in handling occlusions and maintaining stable trajectories as the arm move towards the target point. The results validate the effectiveness of integrating vision-based tracking with kinematic control for precision agricultural tasks. Future work will focus on real-world implementation and the integration of 3D reconstruction techniques for enhanced adaptability in dynamic environments.
@@
thanks: submitted to IFAC AgriControl 2025
@@ 1 Introduction
Despite significant progress in many areas of agricultural robotics (e.g., harvesting, thinning, and crop scouting), pruning still remains a formidable challenge. First, fruit trees exhibit highly unstructured, dynamic environments with complex branch geometries and frequent occlusions by other branches or orchard infrastructure. Second, precise manipulation is required to position a cutting tool at the exact pruning point and orientation while avoiding collisions and excessive forces. Therefore, there is a pressing need for accurate perception, robust point tracking, and agile control schemes to enhance the capability and practical adaptability of robotic pruning systems.
@@
Figure 1: Simulation environment showing the Gazebo setup with the dormant apple trees and a Warthog bot with a mounted UR5e arm.
@@
Automated pruning systems require advanced sensing, modeling, and control to function effectively in dynamic orchard environments. The process begins with accurately sensing and modeling tree structures using sensors like stereo vision or LiDAR, combined with 3D point-cloud registration, to capture branch geometry and spatial arrangement (Botterill et al., 2017; Tabb and Medeiros, 2017). Our previous research demonstrated that accurate branch diameter estimation (Ahmed et al., 2025), along with bud detection and counting (Ahmed et al., 2024), provides essential data for determining optimal crop loads. These measurements inform data-driven pruning rules to identify branches and pruning locations while balancing tree structure maintenance and fruit production. After selecting pruning points, a precise, collision-free trajectory must be planned to guide the cutting tool while avoiding obstacles such as trunks, wires, and posts. However, traditional open-loop or purely position-based methods can accumulate significant errors in real-world conditions, reducing accuracy and precision (You et al., 2020, 2023).
@@
Recent advancements in visual servoing address operational challenges in robotic pruning through closed-loop control based on camera feedback (Shamshiri et al., 2023; Dong and Zhu, 2015). While position-based visual servoing (PBVS) relies on 3D pose estimation, image-based approaches (IBVS) minimize 2D feature errors, offering greater robustness in unstructured orchards where branches, foliage, and infrastructure complicate calibration. Due to its accuracy and precision, visual servoing has been widely explored for tool alignment in dynamic pruning environments. Yandun et al. (Yandun et al., 2021) used deep reinforcement learning (DRL) trained on 3D vine models to navigate cluttered canopies, demonstrating adaptability but requiring extensive training. You et al. (2022) improved this by introducing a hybrid vision/force control framework, where vision-based policies trained on synthetic data guided a cutter to sub-centimeter accuracy, with force feedback mitigating excessive contact with rigid branches. Gebrayel et al. (2024) refined PBVS using iterative closest point (ICP) variants for real-time vine alignment but struggled in highly occluded scenes due to reliance on continuous point-cloud tracking. While RL enables collision avoidance, its dependence on simulated training limits robustness to occlusions, and wind-induced vine motion remains a challenge. These limitations highlight the need for robust visual tracking methods that maintain accuracy despite dense foliage and occlusions.
@@
To overcome orchard challenges such as occlusions, wind-induced movements (Spatz and Theckes, 2013), and variable lighting, tracking methods have advanced beyond keypoint detection and optical flow. These newer approaches handle repetitive textures, occlusions, and rapid motion shifts common in agriculture (Matos et al., 2024). Recent developments include LoCoTrack (Cho et al., 2024), which enhances robustness with 4D correlation volumes for arbitrary point tracking. Tapir (Doersch et al., 2023) employs transformer-based global matching for long-term tracking but remains computationally demanding and struggles in ambiguous regions. CoTracker3 (Karaev et al., 2024a) addresses these limitations with a streamlined architecture that replaces heavy correlation processing with lightweight MLPs, achieving 27% faster performance than LoCoTrack while maintaining accuracy. Its pseudo-labeling approach allows training on real-world videos without manual annotations, improving generalization and reducing data requirements. These features make CoTracker3 well-suited for dynamic orchard environments requiring both precision and efficiency.
This research aims to develop an autonomous robotic system for apple orchard pruning. The robot navigates orchard rows independently, analyzing dormant trees to assess branch geometry, orientation, and bud distribution. Using this structural data, it identifies optimal pruning locations, tracks cutting points with a transformer-based model, and executes precise cuts with a UR5e robotic arm and a cutting tool. To advance autonomous pruning, this work presents an adaptive system that integrates CoTracker3’s robust point-tracking with differential inverse kinematics (IK) and proportional velocity control. Operating primarily in the 2D image plane, the system continuously updates motion based on feature positions, ensuring precise and responsive control. The approach is validated in a Gazebo simulation using a UR5e robotic arm (Universal Robots, Denmark) and an Intel RealSense D435 camera (Intel, California, US).
@@ 2 Methodology
Figure 2: System Architecture (Karaev et al., 2024a): The tracking system uses a camera mounted on the robotic arm to capture visual data and feed to Cotracker3’s perception pipeline to extract convolutional features from each frame. The system analyzes feature correlations between frames to track the pruning point. A transformer iteratively refines the pruning point’s track, confidence, and visibility using previous estimates for accurate tracking.
This section outlines our approach to precision pruning of apple trees. We use the Gazebo simulation platform, integrated with the Robot Operating System (ROS), to create a realistic testing environment, as shown in Figure 1. The robotic system consists of a Universal Robots UR5e manipulator with six degrees of freedom and an Intel RealSense D435 camera mounted on its end-effector. The system’s primary goal is to visually servo the robotic arm toward the pruning branch, ensuring precise end-effector alignment with the target pruning point.
Gazebo serves as the core simulation environment, providing a physics-based framework to model apple trees, the robotic manipulator, and the surrounding workspace. It enables accurate rendering of object interactions, collision dynamics, and sensor feedback, closely replicating real-world pruning conditions. Widely used in robotics research, Gazebo offers a high-fidelity physics engine, sensor integration, and seamless compatibility with ROS, facilitating efficient testing before real-world deployment (Koenig and Howard, 2004).
The vision controller utilizes CoTracker3 to track pruning points across video frames. Currently, pruning points are selected manually, but ongoing research explores AI-driven methods for automatic pruning point detection based on tree geometry. These advancements will be integrated into the system in future iterations. By operating directly in image space rather than reconstructing 3D depth data, the system maintains a streamlined visual servoing approach. The following sections detail the key components of our methodology, including control architecture, feature tracking, controller implementation, and performance evaluation.
@@ 2.1 CoTracker3
CoTracker3 is a transformer-based point tracking model that builds on previous work in CoTracker while simplifying the architecture and improving tracking performance (Karaev et al., 2024b). It is designed for robust tracking of 2D points in video sequences. Unlike traditional methods that track points independently, CoTracker3 leverages joint attention mechanisms to track multiple points jointly, enabling superior tracking performance even under challenging conditions such as occlusions or when points move out of the field of view. This joint tracking approach allows the model to infer the positions of occluded points using information from visible points, significantly improving tracking robustness. By operating on overlapping sliding windows of image frames, CoTracker3 ensures continuity and accuracy over long video sequences.
The CoTracker3 model begins with the user selecting an initial point of interest $(x_{0},y_{0})$ in the camera frame, which serves as the query embedding $\mathbf{q}$. This query embedding initializes the tracking process. CoTracker3 updates the position of the point dynamically across frames by attending to temporal and spatial features in the sliding window of video frames. The tracking operation is formally described as:
$(x_{t},y_{t})=\mathcal{T}\left((x_{t-1},y_{t-1}),\mathbf{F}_{t-1:t},\mathbf{q}\right),$ (1)
where $\mathcal{T}$ represents the CoTracker3 function, $\mathbf{F}_{t-1:t}$ encapsulates the feature embeddings from the sliding window of frames $t-1$ to $t$, and $\mathbf{q}$ is the query embedding derived from the user-selected point $(x_{0},y_{0})$. The model refines the position $(x_{t},y_{t})$ iteratively by leveraging attention mechanisms that link the spatial and temporal contexts. Branch occlusion is a significant challenge in tree pruning scenarios due to the cluttered environment. When a tracked point becomes occluded by branches, CoTracker’s joint attention capabilities infer the most probable position based on the contextual information from visible points and prior frames. This mechanism allows the model to maintain continuity in tracking until the point reappears. To evaluate the performance of the tracking, CoTracker3 computes an error vector between the tracked point $(x_{t},y_{t})$ and the center of the image frame $(x_{c},y_{c})$:
$\mathbf{e}_{t}=\begin{bmatrix}e_{x}\\ e_{y}\end{bmatrix}=\begin{bmatrix}x_{c}-x_{t}\\ y_{c}-y_{t}\end{bmatrix}.$ (2)
The error vector $\mathbf{e}_{t}$ is used to guide the robotic controller in minimizing positional deviations by adjusting the manipulator’s position accordingly. The error vector’s magnitude, $\|\mathbf{e}_{t}\|$, serves as a key metric for assessing tracking accuracy and guiding corrective actions. By combining these mechanisms, CoTracker3 provides a robust and efficient solution for maintaining the pruning point’s position in the camera frame, ensuring seamless and accurate robotic manipulation even in complex environments with branch occlusions.
2.2 Low Level Joint Controller
We adopt a proportional (P) control strategy that operates directly on the 2D image-space error provided by the vision module (CoTracker3). Let $(x_{c},y_{c})$ denote the image center and $(x_{p},y_{p})$ the coordinates of the pruning point as tracked in the camera feed. The error vector is defined as
$\mathbf{e}\;=\;\begin{bmatrix}e_{x}\\ e_{y}\end{bmatrix}\;=\;\begin{bmatrix}x_{c}-x_{p}\\ y_{c}-y_{p}\end{bmatrix}.$ (3)
Denoting proportional gains by $K_{p}^{x}$ and $K_{p}^{y}$, the incremental motion commands in the image plane are calculated as
$\Delta x\;=\;K_{p}^{x}\cdot e_{x},\quad\Delta y\;=\;K_{p}^{y}\cdot e_{y}.$ (4)
@@ 2.2 Low Level Joint Controller
These planar commands define how the end-effector should move in the camera frame to reduce the 2D error. However, the manipulator itself must execute these displacements through joint-space motion. The next subsection details how we map these commands to corresponding joint angle updates via a differential inverse kinematics (IK) approach.
2.3 Differential Inverse Kinematics
To translate the image-based control signals into robot joint motions, we employ a differential inverse kinematics (IK) approach augmented by a deliberate forward step size along the end-effector $z$-axis. This design allows the robot to continuously advance towards the branch while correcting its $x$$y$ alignment based on CoTracker3 feedback.
@@ 2.3 Differential Inverse Kinematics
First, we define the manipulator’s kinematics using the Denavit–Hartenberg (DH) convention. The forward kinematics map joint angles $\mathbf{q}\in\mathbb{R}^{6}$ to a transformation matrix $\mathbf{T}_{\mathrm{end}}(\mathbf{q})\in\mathbb{R}^{4\times 4}$, which encodes the end-effector pose (position and orientation) in 3D space. Denoting by $\mathbf{J}(\mathbf{q})\in\mathbb{R}^{6\times 6}$ the Jacobian of $\mathbf{T}_{\mathrm{end}}$ w.r.t. the joint angles, we have
$\dot{\mathbf{x}}_{\mathrm{end}}\;=\;\mathbf{J}(\mathbf{q})\,\dot{\mathbf{q}},$ (5)
where $\dot{\mathbf{x}}_{\mathrm{end}}\in\mathbb{R}^{6}$ is the spatial velocity of the end-effector (three components for translation, three for rotation), and $\dot{\mathbf{q}}\in\mathbb{R}^{6}$ is the vector of joint velocities.
At each control loop, the vision module (CoTracker3 plus a proportional controller) supplies horizontal and vertical offsets $(\Delta x,\Delta y)$ in image space. We map these offsets to a small planar motion command in the manipulator frame while also prescribing a constant incremental step $\Delta z>0$ along the end-effector’s $z$-axis to move the robot steadily forward toward the pruning target. Symbolically, we assemble a 6D velocity vector:
$\dot{\mathbf{x}}_{\mathrm{end}}\;=\;\begin{bmatrix}\Delta x,\Delta y,\Delta z,0,0,0\end{bmatrix}^{\top},$ (6)
where $\Delta x$ and $\Delta y$ reflect the image-space alignment corrections (appropriately scaled into physical units), and $\Delta z$ is the chosen forward step size in meters per control step (remaining zeros in the equation are for the rotational components).
To realize these Cartesian velocities via joint-space commands, we invert the Jacobian:
$\dot{\mathbf{q}}\;=\;\mathbf{J}^{+}(\mathbf{q})\,\dot{\mathbf{x}}_{\mathrm{end}},$ (7)
where $\mathbf{J}^{+}$ denotes the pseudo-inverse of the Jacobian. Integrating $\dot{\mathbf{q}}$ over a short timestep yields the new joint angles $\mathbf{q}\leftarrow\mathbf{q}+\dot{\mathbf{q}}\,\Delta t$. This iterative scheme continues until the image-space error $(x_{c}-x_{p},\,y_{c}-y_{p})$ remains within a desired tolerance, indicating that the end-effector stays visually locked on the pruning point while advancing forward.
In practice, gains can be tuned to ensure smooth motion, and the forward step size $\Delta z$ can be adjusted according to task needs (e.g., slow approach for delicate cuts). By combining vision-based $x$$y$ corrections with a steady $z$-axis motion in the differential IK framework, the manipulator effectively zeros out lateral errors while proceeding directly toward the pruning location. This integrated control approach yields robust pruning trajectories even under moderate occlusions or shifting tree branches due to wind.

### v2

# An Integrated Visual Servoing Framework for Precise Robotic Pruning Operations in Modern Commercial Orchard

@@ Abstract
This study presents a vision-guided robotic control system for automated fruit tree pruning applications. Traditional pruning practices are labor-intensive and limit agricultural efficiency and scalability, highlighting the need for advanced automation. A key challenge is the precise, robust positioning of the cutting tool in complex orchard environments, where dense branches and occlusions make target access difficult. To address this, an Intel RealSense D435 camera is mounted on the flange of a UR5e robotic arm and CoTracker3, a transformer-based point tracker, is utilized for visual servoing control that centers tracked points in the camera view. The system integrates proportional control with iterative inverse kinematics to achieve precise end-effector positioning. The system was validated in Gazebo simulation, achieving a 77.77% success rate within 5mm positional tolerance and 100% success rate within 10mm tolerance, with a mean end-effector error of 4.28 ± 1.36 mm. The vision controller demonstrated robust performance across diverse target positions within the pixel workspace. The results validate the effectiveness of integrating vision-based tracking with kinematic control for precision agricultural tasks. Future work will focus on real-world implementation and the integration of force sensing for actual cutting operations.
@@
thanks: © 2025 the authors. This work has been accepted to IFAC for publication under a Creative Commons Licence CC-BY-NC-ND.
@@ 1 Introduction
Despite significant progress in many areas of agricultural robotics (e.g., harvesting, thinning, and crop scouting), pruning still remains a formidable challenge. First, fruit trees exhibit highly unstructured, dynamic environments with complex branch geometries and frequent occlusions by other branches or orchard infrastructure. Second, precise manipulation is required to position the end-effector at the correct pruning poses while avoiding collisions and excessive contact forces. Therefore, there is a pressing need for accurate perception, robust point tracking, and agile control schemes to enhance the capability and practical adaptability of robotic pruning systems.
@@
Figure 1: Simulation environment showing the Gazebo setup with the dormant trees and a Warthog bot with a mounted UR5e arm.
@@
Automated pruning systems require advanced sensing, modeling, and control to function effectively in dynamic orchard environments. Previous research demonstrated that accurate branch diameter estimation (Ahmed et al., 2025), along with bud detection and counting (Ahmed et al., 2024), provides essential data for determining optimal crop loads. These measurements inform data-driven pruning rules to identify branches and pruning locations while balancing tree structure maintenance and fruit production. After selecting pruning points, a precise, collision-free trajectory must be planned to guide the end-effector while avoiding obstacles such as trunks, wires, and posts. However, traditional open-loop or purely position-based methods can accumulate significant errors in real-world conditions, reducing accuracy and precision (You et al., 2020, 2023).
@@
Recent advancements in visual servoing address operational challenges in robotic pruning through closed-loop control based on camera feedback (Shamshiri et al., 2023; Dong and Zhu, 2015). While position-based visual servoing (PBVS) relies on 3D pose estimation, image-based approaches (IBVS) minimize 2D feature errors, offering greater robustness in unstructured orchards where branches, foliage, and infrastructure complicate calibration. Yandun et al. (Yandun et al., 2021) used deep reinforcement learning (DRL) trained on 3D vine models to navigate cluttered canopies, demonstrating adaptability but requiring extensive training. You et al. (2022) improved this by introducing a hybrid vision/force control framework, where vision-based policies guided the end-effector to sub-centimeter accuracy, with force feedback mitigating excessive contact with rigid branches. Gebrayel et al. (2024) refined PBVS using iterative closest point (ICP) variants for real-time vine alignment but struggled in highly occluded scenes. These limitations highlight the need for robust visual tracking methods that maintain accuracy despite dense foliage and occlusions.
@@
To overcome orchard challenges such as occlusions, wind-induced movements (Spatz and Theckes, 2013), and variable lighting, tracking methods have advanced beyond keypoint detection and optical flow. CoTracker3 (Karaev et al., 2024a) addresses these limitations with a streamlined architecture that replaces heavy correlation processing with lightweight MLPs, achieving 27% faster performance than LoCoTrack while maintaining accuracy. Its pseudo-labeling approach allows training on real-world videos without manual annotations, improving generalization and reducing data requirements. These features make CoTracker3 well-suited for dynamic orchard environments requiring both precision and efficiency.
This research contributes to the development of an autonomous robotic system for orchard pruning in tree fruit crops (e.g., apple and cherry), with a focus on the vision-based tracking and control strategy required to reach pruning points in complex scenes. While full navigation, structural analysis and collision avoidance are assumed to be externally provided or preprocessed, this work addresses how to effectively track predetermined pruning points and control the robotic arm to achieve precise positioning. The system employs a transformer-based model to robustly track designated pruning points and guides a UR5e robotic arm to these locations. To advance autonomous pruning, this work presents a control strategy that integrates CoTracker3’s robust point-tracking with iterative inverse kinematics and proportional control. The controller focuses on reaching target points using visual feedback, with depth monitoring providing a simple stopping mechanism when approaching within 20 cm of the target. Operating primarily in the 2D image plane, the system continuously updates motion based on tracked point positions, ensuring precise and responsive control. The approach is validated in a Gazebo simulation using a UR5e robotic arm (Universal Robots, Denmark) and an Intel RealSense D435 camera (Intel, California, US).
@@ 2 Methodology
Figure 2: System Architecture: The tracking system uses a camera mounted on the robotic arm to capture visual data and feed to CoTracker3’s perception pipeline (Karaev et al., 2024a) to extract convolutional features from each frame. The system analyzes feature correlations between frames to track the pruning point. A transformer iteratively refines the pruning point’s track, confidence, and visibility using previous estimates for accurate tracking.
This section outlines this study’s approach to precision pruning of orchard trees. The Gazebo simulation platform, integrated with the Robot Operating System (ROS) is used to create a controlled testing environment, as shown in Figure 1. The robotic system consists of a Universal Robots UR5e manipulator mounted on a Clearpath Warthog mobile platform, with an Intel RealSense D435 camera attached to the manipulator’s flange. The system’s primary goal is to visually servo an end-effector toward a target branch, ensuring precise end-effector alignment with the selected pruning point.
Gazebo serves as the core simulation environment, providing a physics-based framework to model orchard trees, the robotic manipulator, mobile platform, and the surrounding workspace. The simulated orchard features dormant trees arranged in a tall spindle architecture, representing high-density planting systems commonly used in modern commercial orchards. Gazebo enables rendering of object interactions and collision dynamics, though the simulated sensor feedback may differ from real-world conditions. The simulation does not fully account for sensor noise, lighting variations, tree architecture differences and environmental complexities present in field conditions. Despite these limitations, Gazebo’s high-fidelity physics engine and seamless compatibility with ROS provide a valuable platform for initial validation before real-world deployment (Koenig and Howard, 2004).
The vision controller utilizes CoTracker3 to track pruning points across video frames. Currently, pruning points are selected manually, but ongoing research explores AI-driven methods for automatic pruning point detection based on tree geometry. These advancements will be integrated into the system in future iterations. The system operates primarily in image space for visual servoing control, with depth information used solely as a stopping mechanism when the robot approaches within 20cm of the tracked pruning point. The following sections detail the key components of this study’s methodology, including control architecture, feature tracking, controller implementation, and performance evaluation.
@@ 2.1 CoTracker3
CoTracker3 is a transformer-based point tracking model (see Figure 2) that leverages joint attention mechanisms to track multiple points simultaneously, enabling robust performance under challenging conditions such as occlusions (Karaev et al., 2024b). Unlike traditional methods that track points independently, CoTracker3 can infer occluded point positions using information from visible points and prior frames, significantly improving tracking robustness in cluttered environments.
The model begins with user selection of an initial point $(x_{0},y_{0})$ which serves as a query embedding for tracking initialization. The system monitors depth and automatically stops visual servoing when the tracked point reaches within 20cm of the camera as a proximity-based termination condition.
For visual servoing control, the system computes an error vector between the tracked point $(x_{t},y_{t})$ and the image center $(x_{c},y_{c})$:
$\mathbf{e}_{t}=\begin{bmatrix}x_{c}-x_{t}\\ y_{c}-y_{t}\end{bmatrix}$ (1)
This error vector drives the visual servoing control law, guiding the robotic manipulator to center the pruning point in the camera view for precise end-effector alignment.
2.2 Image-Based Visual Servoing Controller
We adopt a proportional (P) control strategy that operates directly on the 2D image-space error provided by CoTracker3. The proportional controller generates incremental motion commands in the image plane according to
$\Delta x\;=\;K_{p}^{x}\cdot e_{x},\quad\Delta z\;=\;K_{p}^{z}\cdot e_{y}.$ (2)
where $K_{p}^{x},$ and $K_{p}^{z}\in\mathbb{R_{>\textrm{0}}}$ are empirically tuned positive proportional gains that map pixel errors to metric displacements in the robot frame. In our notation, $\mathbb{R_{>\textrm{0}}}$ denotes a positive real number. Horizontal image errors ($e_{x}$) map to lateral end-effector motions ($\Delta x$), while vertical image errors ($e_{y}$) map to vertical end-effector motions ($\Delta z$) as established by the camera mounting configuration. These conservative gain values ensure stable convergence without overshoot while maintaining responsive tracking performance. The error terms $(x_{c}-x_{t})$ and $(y_{c}-y_{t})$ naturally encode the required motion direction: when the pruning point appears to the right of the image center ($x_{t}>x_{c}$), the negative error drives the end-effector leftwards to recenter the target. Additionally, the system applies a constant forward motion ($\Delta y$) increment per control cycle to steadily advance the end-effector toward the pruning target.
@@ 2.2 Image-Based Visual Servoing Controller
These Cartesian commands define how the end-effector should move in the camera frame to reduce the 2D tracking error. However, the manipulator must execute these displacements through joint-space motion. The next subsection details how these Cartesian commands are mapped to corresponding joint angle updates via a iterative inverse kinematics approach.
2.3 Iterative Inverse Kinematics
To translate the image-based control signals into robot joint motions, an iterative inverse kinematics (IK) approach with singularity handling and joint limit enforcement is employed. The approach iteratively solves for pose errors and updates joint angles until convergence.
@@ 2.3 Iterative Inverse Kinematics
The manipulator’s kinematics are defined using the Denavit-Hartenberg (DH) convention with parameters specific to the UR5e robot. Forward kinematics map joint angles $\mathbf{q}\in\mathbb{R}^{6}$ to end-effector poses in 3D space via a transformation matrix $\mathbf{T}_{\mathrm{end}}(\mathbf{q})\in\mathbb{R}^{4\times 4}$, which encodes the end-effector pose in 3D space.
At each control iteration, the desired end-effector pose is computed by applying the visual servoing corrections to the current pose. The desired position incorporates the lateral correction $\Delta x$, forward motion $\Delta y$ and vertical correction $\Delta z$ while maintaining the current orientation to preserve end-effector alignment.
The IK solver iteratively minimizes the pose error between current and desired poses:
$\mathbf{e}_{\text{pose}}=\begin{bmatrix}\mathbf{e}_{\text{pos}}\\ \mathbf{e}_{\text{orient}}\end{bmatrix}=\begin{bmatrix}\mathbf{p}_{\text{target}}-\mathbf{p}_{\text{current}}\\ \text{Log}_{SO(3)}\left(\mathbf{R}_{\text{target}}\mathbf{R}_{\text{current}}^{T}\right)\end{bmatrix}$ (3)
where $\mathbf{e}_{\text{pos}},\mathbf{e}_{\text{orient}}\in\mathbb{R}^{3}$ are the position and orientation errors respectively, and $\mathbf{R}_{\text{target}},\mathbf{R}_{\text{current}}\in\mathbb{R}^{3\times 3}$ are the target and current rotation matrices.
To address singularities and ensure numerical stability, we employ a damped pseudo-inverse formulation:
$\Delta\mathbf{q}=\mathbf{J}^{T}(\mathbf{J}\mathbf{J}^{T}+\lambda\mathbf{I})^{-1}\mathbf{e}_{\text{pose}}$ (4)
where $\mathbf{J}\in\mathbb{R}^{6\times 6}$ is the numerical Jacobian computed using finite differences, $\lambda=10^{-4}$ is the damping factor, and $\mathbf{I}\in\mathbb{R}^{6\times 6}$ is the identity matrix.
The algorithm incorporates error-dependent step sizing to regulate convergence:
$\alpha=\text{clip}\left(0.01\cdot(1.0+\|\mathbf{e}_{\text{pose}}\|),0.001,0.05\right)$ (5)
with step size reduction near singularities when $\det(\mathbf{J}\mathbf{J}^{T})<10^{-6}$. The joint angles are updated as $\mathbf{q}\leftarrow\mathbf{q}+\alpha\cdot\Delta\mathbf{q}$, with step magnitude limited to 0.1 radians per iteration to prevent discontinuous motion.
The iteration terminates when both position and orientation errors satisfy convergence criteria ($\|\mathbf{e}_{\text{pos}}\|<10^{-4}$ m and $\|\mathbf{e}_{\text{orient}}\|<10^{-3}$ rad) or after a maximum of 1000 iterations. This iterative formulation ensures stable convergence while enforcing joint limits and preventing excessive rotations, particularly for wrist joints where angular displacements are constrained to minimize unnecessary motion. The visual servoing process continues until depth monitoring detects that the tracked point is within 20cm of the camera, at which point, motion commands cease to prevent contact with the pruning target. Currently, the system relies solely on this depth-based termination criterion. Future field work will incorporate force sensing control for more sophisticated contact detection and control.